OpenStudy (friendzone):

A two-sided coin (heads/tails) is flipped once, and one card is drawn from a standard 52 deck of playing cards. Calculate the probability of each given situation, and drag the correct answer to it. Flipping a Heads AND drawing a Heart Flipping a Heads OR drawing a Heart Flipping a Heads Drawing a Heart NOT drawing a King Flipping a Tails AND drawing a King Flipping a Tails OR drawing a King Drawing a King OR drawing a Heart Flipping a Heads OR flipping a Tails Drawing a King AND drawing a Heart BANK: Move to Top 1 1 26 1 2 1 4 1 52 1 8 12 13 15 26 3 4 4 13

OpenStudy (tkhunny):

Are they independent events? That makes it easier. Just multiply the two probabilities. P(Flipping a Heads) = 1/2 AND P(drawing a Heart) = 1/4 Finish that one and show another.

OpenStudy (friendzone):

thanks it just keeps going

OpenStudy (xapproachesinfinity):

if they two independent events (problem should state it) P(flipping a head or drawing a heart card)=P(Head)+P(Heart Card)-P(Heads)P(Heart Card)